You have found the following stock quote for RJW Enterprises, Inc., in the financial pages of today’s newspaper.

52-WEEK YLD VOL NET
HI LO STOCK (DIV) % PE 100s CLOSE CHG
130.25 88.38 RJW 2.50 2.0 15 17,907 ?? .73

What was the closing price for this stock that appeared in yesterday’s paper? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Yesterday’s closing price $ 125

^CORRECT

If the company currently has 28 million shares of stock outstanding, what was net income for the most recent four quarters? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Net income $   

^QUESTION
Answer #1

Solution:-

Today's paper would contain yesterday's closing prices and yesterday's paper would contain closing prices for day before yesterday. The question requires us to calculate the closing price as of yesterday's paper, i.e. the closing price of the security for day before yesterday

We are given that the stock dividend was 2.5 and dividend yield is 2%. The dividend yields are always calculated based on latest price. Therefore, the dividend yield of 2% is the yield based on yesterday's closing price as appearing in today's paper.

Dividend yield= 2%

Dividend/yesterday's closing price= 2%

2.5/Yesterday's closing price= 2%

Yesterday's closing price= $125

Closing price of day before yesterday (i.e. closing price as appeared in yesterday's paper)= 125-change = 125-(-0.73) = $125.73

Total market cap= 125*28m= $3,500 million

PE ratio= 15 times

Net income for last year= 3,500/15= $233.3 million